OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (anonymous):

y = 1/3x + 14/3 y = 3x + 10 y = -3x - 2 y = -1/3x + 10/3

OpenStudy (anonymous):

i say A

Nnesha (nnesha):

do you know what is the slope of this equation y = -3x +6 ??

OpenStudy (anonymous):

wait no that would be perpindicular

Nnesha (nnesha):

^^^

OpenStudy (anonymous):

-3 is the slope

OpenStudy (anonymous):

so it would be C?

OpenStudy (anonymous):

Bcause it has to have the same slope to be parrelel?

Nnesha (nnesha):

yes so u need to find equation which suppose to be parallel line so slope of new equation suppose to be what ??

Nnesha (nnesha):

yep you are lucky there is only one answer choice with same slope ;)

OpenStudy (anonymous):

haha ok thanks

Nnesha (nnesha):

np :)

OpenStudy (anonymous):

i got cpnfused with perpendicular and parrelel

Nnesha (nnesha):

yep parllel lines have same slope but perpendicular line should be negative reciprocal

OpenStudy (anonymous):

yup ok thanks again

OpenStudy (anonymous):

if i need help again can you help?

Nnesha (nnesha):

i'll try :)

OpenStudy (anonymous):

y = m x + b {x = -2, y = 4, m = -3} b = 2

Nnesha (nnesha):

b is not 2

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @robtobey y = m x + b {x = -2, y = 4, m = -3} b = 2 \(\color{blue}{\text{End of Quote}}\) 4 = -2(-3) + b 4 = 6 + b subtract 6 both side -2 = b i guess there is a typo which is 2 instead of -2